OpenStudy (anonymous):

Solve the systems using substitution. 6x + 2y = 0 4x + y = -1

OpenStudy (lynfran):

frm equation 2, let y=-1-4x now sub that into equation 1

OpenStudy (boldjon):

for number 1 Let's solve for x. 6x+2y=0 Step 1: Add -2y to both sides. 6x+2y+−2y=0+−2y 6x=−2y

OpenStudy (boldjon):

Step 2: Divide both sides by 6. 6x/6=-2y/6

OpenStudy (boldjon):

so u'll get something like -1/3

OpenStudy (boldjon):

check to make sure though

OpenStudy (anonymous):

In Mathematica, the equations can can be evaluated at x and y equals -1 and 3 respectively. If the LHS and RHS for each equation, post substitution, are equal, then "True" is returned, else "False". /. is the replacement operator. { 6x+2y=0, 4x+y=−1 } /. {x→−1,y→3} {True, True}